OpenStudy (anonymous):

To approach a runway, a plane must begin a 7 degree descent starting from a height of 3 miles above the ground. To the nearest mile, how many miles from the runway is the airplane a the start of this approach?

OpenStudy (anonymous):



OpenStudy (anonymous):

cot 7=cot(90-83)=tan 83

OpenStudy (anonymous):

x=3 tan 83=3*8.14=24.42 mile or

OpenStudy (koikkara):

Approx , \(X\)= 11.64rad ?